‎6[sup]x-1[/sup] =3[sup]2x[/sup]
1
Expert's answer
2011-05-17T09:34:58-0400
3(x-1)* 2(x-1) = 32x
2(x-1)= 3(2x-x+1)
2(x-1)= 3(x+1)
2(x-1)= 9*3(x-1)
(2/3)(x-1)= 9
x-1 =log(2/3)9
x=1+log(2/3)9=log(2/3)2/3* 9=log(2/3)6